PS-IS-126/UG

 

MINUTES OF THE PS-ISOLDE COORDINATION MEETING held on 15 October 2002

 

 

Present:

P. Butler, R. Catherall, A. Dorsival, V. Fedosseev, G. Focker, U. Georg, T.Giles, U. Koester, G. Le Scornet, J. Lettry, M. Lindroos, S. de Man, S. Meunier, T. Nilsson, J. Parra-Lopez, I. Podadera Aliseda, F. Wenander

 

ISOLDE at URL: http://isolde.cern.ch

Minutes at URL: http://ps-div.web.cern.ch/ps-div/Meetings/ISOLDE/TechMtgs.html

Consolidation projects at URL: http://isolde-consolidation.web.cern.ch/isolde-consolidation/

PSMB meetings at URL: http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html

1. HRS:

- The target was changed from UC2 to Nb on Thursday during the MD.

- The camera used inside the beam tunnel to observe the HRS target shelves is broken.

- Scanner SC260 is broken (movement error)

- There is an alignment mismatch between the slit behind the 2nd separator magnet and the fixed needle beam scanner.

2. GPS:

- The run with the Ta-foil target went well and the production of Be and Li was quite stable.

The run was interrupted by the following events:

i) Tuesday, the GPS transformer switched off (reason unknown)

ii) Thursday, after the HRS target change the concrete shielding door in the proton beam tunnel was closed but failed to send the correct signal to the Booster for proton delivery.

iii) Saturday, a 24V power supply at the PLC for target heating broke and the heating stopped.

To continue the run the target and line supplies were switched to manual mode.

3. Tape station

- The tape station needs still a major check-up. Currently it is back usable but without the optocoupler stopping it automatically once the end of the tape is reached.

4. RILIS

- Lasers are set now for Zn ionisation. Two new crystals didn't perform well so we are currently running without spare.

 

5. Miscellaneous

- Mats is writing the ISOLDE chapter for the PS divisional report, please send your input to him.

- A noise level measurement has been performed in the ISOLDE experimental hall and the control room. While the level in the hall (68 dB/A) is within recommended levels for such environments the noise inside the control room (62 dB/A at the desks) is above the recommendations (60dB/A) for such rooms.

It can be reduced to 59 dB/A by keeping doors and racks closed. A big contribution is apparently coming from fans in NIM crates which might not all be necessary.

Next meeting:

Tuesday 22 October 2002 at 11.00 in the EP Conference Room A, building 304-1-001.

 

Minutes by Uwe Georg